OpenStudy (anonymous):

Express the complex number in trigonometric form -3i

OpenStudy (alekos):

-3i = 0 - 3i if a complex number has the form a+bi and you want to convert to rcisθ \[r =\sqrt{a^{2}+b^{2}}\] \[\theta = \arctan b/a\]

OpenStudy (anonymous):

thank you!